OpenStudy (anonymous):

. Which ordered pair in the form (a, b) is a solution of this equation? 4a – 5b = 24 A. (0, 6) B. (6, 0) C. (–5, –2) D. (–2, –5)

OpenStudy (zale101):

plug in the ordered pairs to the equation and see if the equation does equal to 24

OpenStudy (anonymous):

4(0)-5(6) = 24 0-30 = 24 no...try next ordered pairs

OpenStudy (anonymous):

um can you explain a bitmore im not strong at theis subject but dont give me the answer

OpenStudy (anonymous):

(6,0)= (a,b) Equation 4a-5b = 24 Put 6 in for a and 0 in for b 4(6)+5(0)=24

OpenStudy (anonymous):

when you simplify & both sides of the equation are equal, u have the solution

OpenStudy (anonymous):

(-5,-2) 4a-5b = 24 4(-5)+5(-2)=24 -20+10 = 24 -10 = 24 no...this ordered pair isn't a solution

OpenStudy (disco619):

So the equation is: \[4a – 5b = 24\] And we have been given a couple of Pairs in the form (a, b) Which are: A. (0, 6) B. (6, 0) C. (–5, –2) D. (–2, –5) So in all of the pairs, the left digit is "a" and the right digit is "b". In (0, 6): a = 0 b = 6 So in the equation: \[4a – 5b = 24\] We would replace the value of "a" in pair which in Option A is "0" with the "a" in equation. We would do the same with the value of "b". \[4(0) – 5(6) = 24\] And then solve it to see if we get 24 \[0 – 30 = 24\] \[-30 = 24\] This would be wrong, hence we would do the same with Option B, C and D. And see which one is right and comes up with the Answer 24. That Option would be correct.

OpenStudy (anonymous):

ohh okay i get it can you help mr on another one

OpenStudy (disco619):

I can try

OpenStudy (anonymous):

okay

OpenStudy (anonymous):